Bob owns a two-stock portfolio that invests in Falcon Freight Company (FF) and Pheasant Pharmaceuticals (PP). Three-quarters of Bob’s portfolio value consists of Falcon Freight’s shares, and the balance consists of Pheasant Pharmaceuticals’s shares. Each stock’s expected return for the next year will depend on forecasted market conditions. The expected returns from the stocks in different market conditions are detailed in the following table: Market Condition Probability of FF PP Occurrence Strong 20% 50% 70% Normal 35% 30% 40% Weak 45% -40% -50% Calculate expected returns for the individual stocks in Bob’s portfolio as well as the expected rate of return of the entire portfolio over the three possible market conditions next year. • The expected rate of return on Falcon Freight’s stock over the next year is . • The expected rate of return on Pheasant Pharmaceuticals’s stock over the next year is . • The expected rate of return on Bob’s portfolio over the next year is .
Answer #1

The expected rate of return on Falcon Freight’s stock over the next year is=0.20*50%+0.35*30%+0.45*(-40%)=2.5000%

The expected rate of return on Pheasant Pharmaceuticals’s stock over the next year is=0.20*70%+0.35*40%+0.45*(-50%)=5.5000%

The expected rate of return on Bob’s portfolio over the next year is=3/4*2.5%+1/4*5.5%=3.2500%